OpenStudy (kj4uts):

A breeding group of beavers is introduced into a protected area. After t years the number of beavers in the area is modeled by the function N(t)= 54/0.35+0.68^t 1. How many beavers were initially introduced? 2. Estimate the number of beavers after 5 years. 3. Determine the change in the beaver population between t = 5 and t = 10. (Note: All answers are whole numbers.)

OpenStudy (kj4uts):

My answers that I have so far are (I am not sure if I did them right): 1. N(t)= 54/0.35+0.68^0 = 40 2. N(t)= 54/0.35+0.68^5 = 109 3. I am not sure?